OpenStudy (anonymous):

Find the quotient of 31,017 ÷ 49 and then show how you would estimate the problem. 62331,000 ÷ 50 = 620 64331,000 ÷ 50 = 620 62330,000 ÷ 50 = 600 63330,000 ÷ 50 = 600 this is confusing me

OpenStudy (triciaal):

what did you get for the quotient ?

OpenStudy (anonymous):

633

OpenStudy (triciaal):

correct

OpenStudy (anonymous):

but what do i do after that do i estimate 633 into 630 and divide that by 49 to get 50?

OpenStudy (triciaal):

do you see 49 is close to 50 if you divide by a higher number the quotient is less do you see

OpenStudy (aum):

Could you post a screenshot of this question?

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

should it be d?

OpenStudy (triciaal):

i think you have the original wrong 31, 017 * 2 = 62,034 but you have thousand, thousand

OpenStudy (aum):

The question and the answer choices are not clear to me. I'd like to see the original question. If it is online, you can take a screenshot and post the original problem. If it is in a book you can take a picture and upload it.

OpenStudy (anonymous):

that is the question and answer but what im trying to ask is how would u estimate to find the answer

OpenStudy (aum):

If I were asked to estimate 31,017 ÷ 49 I would say 31,017 ÷ 49 is approximately equal to 31,000 ÷ 50 = 2 * 31,000 / 100 = 2 * 310 = 620 (estimate). But the answer choices does not make sense to me.

OpenStudy (anonymous):

same with me i dont get them

OpenStudy (aum):

Can you just post a pic of the question just to see if there is something you are overlooking?

OpenStudy (triciaal):

did you check the original?

OpenStudy (anonymous):

this is the original

OpenStudy (triciaal):

well I really don't know. " Find the quotient of 31,017 ÷ 49 and then show how you would estimate the problem." It appears like the problem to estimate is not even there then the rest is just the choices for what a person would get when the estimate is figured.

OpenStudy (anonymous):

well thanks for trying but this is soooo confusing

OpenStudy (triciaal):

dividing by 49 is close to dividing by 50 so that part is very clear the number to divide I thought it would be 2 * 31017 but the number you have is 62034000. I really still think the problem to estimate is actually not there. sorry

OpenStudy (anonymous):

633